# ProST: Progressive Sub-task Training for Pareto-Optimal Multi-agent Systems Using Small Language Models

### v1

@@ 7 Analysis
Table 2 shows that random selection yields the worst performance, whereas progressively learning more subtasks either via (Decrement or Ours) provides gains over training on all subtasks from the start. However, Ours, the progressive curriculum whereby we learn the task-specific subtasks first, is still substantially better than its inverse Decrement, likely because it better reflects the natural dependencies between the subtasks. However, we see that SGC scores are the same for All, Decrement, and Ours, showing that selection does not affect consistently solving tasks across the same scenarios.
ProST reduces error rates in subtasks
@@
We compare the trajectory lengths of single and multi-agent systems. Figure 6 shows the distribution of tokens for the 14B based systems in successfully completed test set tasks compared to the unsuccessful ones. The trajectory lengths, i.e, the average number of tokens, for unsuccessful tasks is higher than for successful ones for all systems with higher variance. Unsuccessful tasks are likely to have more trial-and-error behavior and more errors, which add to the trajectory length. Multi-agent systems have longer trajectories for successful tasks compared to single-agent systems in part because of the overheads involved in separating the roles and the resultant communication. Fine-tuning, standard, and ProST have shorter trajectories for successful tasks, likely due to reduced subtask error rates.
@@
Table 3: TGC and SGC scores on AppWorld Test-Normal benchmark for different baselines and models (Llama-3.1-8B and Phi-4-14B). We notice similar performance tradition between Qwen series models and Llama-3.1-8B & Phi-4-14B. We observe that multi-agent systems trained with ProST gives best performance in all settings. SA = Single Agent system, MA = Multi-Agent system, FT = Standard finetune, ProST = Finetune with Progressive Subtask Training.
@@
AppWorld also provides an out-of-distribution Test-Challenge test set, including tasks that use at least one unseen application. We evaluate Qwen-2.5-Coder-14B and Phi-4-14B variants on this test set (see Table 4). Qwen-2.5-Coder-14B [FT] Single Agent (SA) performs better than the Multi-Agent (MA) baselines. Manual analysis shows this lack of generalization stems from ineffective exploration of the context and APIs associated with the new applications in this test set. For instance, in the task How many hours does the battery of the ecobee SmartCamera last? Please answer as per its Amazon reviews or questions/answers and only trust information from its verified purchasers, MA only explores ‘reviews’ for the solution, ignoring ‘questions/answers’. Nevertheless, MA [ProST] is still better than MA [FT], showing that Progressive Subtask Training is still better than standard fine-tuning for multi-agent settings. On the other hand, Phi-4-14B shows consistent performance increase on the Test Challenge set similar to the Test Normal set. Phi-4-14B MA [ProST] outperformed both Qwen-2.5-Coder-14B SA [FT] and Phi-4-14B SA [FT] and achieves the best performance $17.8\%$, and $8.6\%$ in both TGC and SGC showing utility in OOD settings. These results indicate that [ProST] can be beneficial, but the overall low performance indicates the challenges in OOD settings for small language model (SLM) based agents.
@@
Table 4: Performance comparison in single-agent and multi-agent settings on Test Challenge set.
@@ 8 Conclusion
We use AppWorld as a test bed for complex reasoning problems. While our design for the agents i.e., their roles, are broadly applicable to many agentic frameworks, and is not AppWorld specific, some aspects of our modeling and agentic design are influenced by specifics of AppWorld. For example, we showed through ablation studies that the order of introducing subtasks in ProST does indeed matter. However, the choice of introducing the first and last subtasks at the end of training is specific to AppWorld, and it may not generalize well to other benchmarks.
@@
The training dataset is LLM generated. While we only use successful trajectories, the overall quality of these trajectories (e.g., diversity, reasoning correctness etc.) has not been evaluated. Also, the instances are created using a closed-source LLM (Claude-3.7-Sonnet). We will release these instances to ensure replicability and comparisons against other methods for training agents. However, the standard issues of using closed-source LLM will make it difficult to compare against other means of generating training data, unless the method is replicated. We will release code and data to support replicability as much as possible.
@@
Since we finetune models on training data with limited scope, models may have learned short-cuts, or overfit to the distribution of the training data. The mixed results on out-of-distribution test data in table 4 suggests scope for future work.
@@ Acknowledgements
This work is supported in part by a SUNY-IBM Artificial Intelligence Collaborative Research Alliance grant, an Amazon Research Award 2023, and a NSF NAIRR compute credits grant #240140.
Appendix A Creating the Training Data
A.1 Single Agent Setting
@@
AppWorld Trivedi et al. (2024) provides training pairs consisting of input prompts and output states. Specifically, it offers 90 training and 57 development tasks, each with its corresponding output state (i.e., Python solution). We use all 147 tasks to create a fine-tuning dataset for both single-agent and multi-agent setups. To increase the dataset size and introduce solution diversity for each task, we use Llama-3.3-70B-Instruct Meta (2025) as the Re-Act agent across 20 different temperature settings. We start at a temperature of $0.05$ and increment by $0.05$ each time, up to $1.0$. To reflect practical scenarios, we first attempt to solve all $147\times 20=2,940$ tasks without providing the output state to the agent. Our base Llama-3.3-70B-Instruct model successfully solves 1,034 out of 2,940 candidate trajectories across the different temperature settings. For the remaining tasks, we provide abstract instructions instead of direct code. We generate these instructions in natural language using Llama-3.3-70B-Instruct. By feeding these abstract instructions to the agent for the 1,906 unsolved candidates, we obtain 1,396 new gold trajectories. For the final 510 candidates, we convert the Python code solutions into pseudocode using Llama-3.3-70B-Instruct. We avoid giving direct Python code as hints because the agent tends to reproduce the code verbatim. Assisting in pseudocode format yields 414 additional gold trajectories. In total, we collect 2,844 gold trajectories. This approach ensures diverse solutions for each task across different temperature settings. When our agent fails to solve tasks in the first phase (i.e., limit exceeds), we use our curated abstract assistance. We did not generate new training tasks for AppWorld, since creating and validating new tasks requires specific environment support, which is challenging to set up.
A.2 Single Agent Trajectory to Multi-agent Trajectory
We ensure that each agent in the multi-agent system is trained effectively to handle its specific role while maintaining overall system coherence. The most challenging part is creating a dataset for the Orchestrator agent. There are two options: first, we can use the base Llama-3.3-70B-Instruct model to act as the Orchestrator agent and generate plans, and use the same Llama-3.3-70B-Instruct model to act as the Executor and Critic agents. This approach has a significant drawback: open-source models such as Llama-3.3-70B-Instruct are not trained on such long-horizon tasks, so the plans are not robust and the overall trajectories are less informative. Instead, it is easier to create single-agent trajectories, and they are more informative. Alternatively, we could use closed-source models such as Claude, Gemini, or OpenAI models, but the cost of doing so is very high. The second, more efficient, and cost-effective approach that we follow is to use Claude 3.7-Sonnet Anthropic (2025) to transform these gold Re-Act trajectories into multi-agent trajectories (See the prompt in Figure 12 in Appendix D). For this transformation, it cost us around 400 USD on the Anthropic API. As Claude 3.7 is a very powerful model, it can generate very rich plans and also divide the actions for each of the plans very effectively. Claude 3.7 does not generate any code; instead, it generates robust plans from the already given gold Re-Act trajectories. We then manually check 50 randomly selected datapoints and find that the generated plans and actions belong to the same subtasks and are correct. We further systematically check the validity of the transformed multi-agent trajectories by running them in the AppWorld environment to see whether the trajectories still solve the task. Moreover, Figure 7 shows the error rates for each subtask of the multi-agent trajectories. It shows our dataset covers error correction patterns across all subtasks. We finetune specialized agents of different sizes (7B and 14B), as well as single agents as baselines. Agents in the multi-agent systems are trained on their respective datasets in two ways: (1) standard supervised finetuning, and (2) Progressive Subtask Training (ProST) (details in Section 4.)
@@ A.2 Single Agent Trajectory to Multi-agent Trajectory
Appendix B Implementation Details
B.1 Training and Evaluation Configuration
Our base models are finetuned with Low-Rank Adaptation (LORA) Hu et al. (2021). We use rank (r=16), alpha ($\alpha$=32) and dropout = 0.05. We use a learning rate of 2.0e-4 with a cosine scheduler. To stabilize training, we set the gradient clipping to 1.0 and the weight decay to 0.01. We train the model for 5 epochs. We use a warm-up ratio of 0.1. The maximum sequence length is set to 20,480 tokens, as our training trajectories are long. All of our experiments are conducted on 4 NVIDIA H100 GPUs with 80GB of memory each. We use Accelerate integrated with DeepSpeed (ZeRO-3) for multi-GPU training. This setup provides efficient memory management and optimization. For both training and evaluation, each agent took  8 GPU hours. For evaluation, we use vLLM Kwon et al. (2023) servers. For both single-agent and multi-agent setups, we perform inference at T=0.1 and T=0.2 for Test Normal set and Test Challenge, respectively. We set the maximum sequence length to 65536 and enable prefix caching. For the single-agent setting, we allow up to 50 turns for evaluation. For the multi-agent, Orchestrator can plan up to 12 subtasks, and each time Executor allows max 15 turns (for the Test Challenge set, we set max 20 turns) to complete each subtask. A task is considered failed if an agent calls the task completion API with a fail status or doesn’t complete the task within the maximum allotted limits.
B.2 Random Subtask Selection Strategy
B.3 ProST Details
Appendix C Appworld Details
AppWorld Trivedi et al. (2024) is a framework designed to assess the ability of autonomous agents in solving real-world problems by interacting with an environment of everyday applications. The framework provides the user with two components: the AppWorld Execution Engine and AppWorld Benchmark.
Appendix D Prompts Details

### v2

@@ 7 Analysis
Figure 5: Failure mode analysis using the MAST cemri2025multiagentllmsystemsfail taxonomy (Specification Issues, Inter-Agent Misalignment, Task Verification) and the LLM-as-a-Judge pipeline (Llama-3.3 70B as the judge). Progressive Subtask Training (ProST) consistently reduces error modes across all categories.
@@
Table 2 shows that random selection yields the worst performance, whereas progressively learning more subtasks either via (Decrement or Ours) provides gains over training on all subtasks from the start. However, Ours, the progressive curriculum whereby we learn the task-specific subtasks first, is still substantially better than its inverse Decrement, likely because it better reflects the natural dependencies between the subtasks. However, we see that SGC scores are the same for All, Decrement, and Ours, showing that selection has no affect on solving tasks across same scenarios.
ProST reduces error rates in subtasks and MAST taxonomy
@@
To analyze the errors further, we adopt MAST (Multi-Agent System Failure Taxonomy) cemri2025multiagentllmsystemsfail, an empirically grounded taxonomy for categorizing errors in multi-agent systems, which categorizes failures into Specification Issues, Inter-Agent Misalignment, and Task Verification. We use their validated LLM-as-a-Judge pipeline (with Llama-3.3 70B as the judge) to classify errors across these categories. As shown in Figure 5, Progressive Subtask Training (ProST) consistently reduces error rates across all categories and subcategories, with highest improvements in inter-agent coordination failures.
@@
We compare the trajectory lengths of single and multi-agent systems. Figure 7 shows the distribution of tokens for the 14B based systems in successfully completed test set tasks compared to the unsuccessful ones (see Figure 9 in the Appendix for the full results across 7B, 14B, and 32B models) . The trajectory lengths, i.e, the average number of tokens, for unsuccessful tasks is higher than for successful ones for all systems with higher variance. Unsuccessful tasks are likely to have more trial-and-error behavior and more errors, which add to the trajectory length. Multi-agent systems have longer trajectories for successful tasks compared to single-agent systems in part because of the overheads involved in separating the roles and the resultant communication. Fine-tuning, standard, and ProST have shorter trajectories for successful tasks, likely due to reduced subtask error rates.
@@
Table 3: TGC and SGC scores on AppWorld Test-Normal benchmark for different baselines and models (Llama-3.1-8B and Phi-4-14B). We notice similar performance tradition between Qwen series models and Llama-3.1-8B & Phi-4-14B. We observe that multi-agent systems trained with ProST have the best performance in all settings. SA = Single Agent system, MA = Multi-Agent system, FT = Standard finetune, ProST = Finetune with Progressive Subtask Training.
@@
AppWorld also provides an out-of-distribution Test-Challenge test set, with tasks requiring more API calls and use at least one unseen application. We evaluate ProST on this test set, presenting results for Qwen-2.5 Coder 14B and Phi-4 14B in Table 4. We find that Phi-4-14B with ProST outperforms other fine-tuning and multi-agent baselines with 17.817.8% in TGC and 8.68.6% in SGC. While Qwen-2.5-Coder-14B’s fine-tuned single-agent achieves the best performance in its model family, its ProST variant has higher accuracy than the corresponding fine-tuned multi-agent setup. These results show that ProST, with appropriate base models, can help improve OOD generalization. More results for this test set are present in Appendix Table 6.
@@
Table 4: Performance comparison in single-agent and multi-agent settings on Test Challenge (Test-C) set for Qwen-2.5-Coder-14B and Phi-4-14B.
@@ 8 Conclusion
We use AppWorld as a test bed for complex reasoning problems. While our design for the agents i.e., their roles, and ProST training paradigm are broadly applicable to many agentic frameworks, and is not AppWorld specific (detailed explanation in Appendix D.1), some aspects of our modeling and agentic design are influenced by specifics of AppWorld. For example, we showed through ablation studies that the order of introducing subtasks in ProST does indeed matter. However, the choice of introducing the first and last subtasks at the end of training is specific to AppWorld, and it may not generalize well to other benchmarks.
@@
The training dataset is LLM generated. While we only use successful trajectories, the overall quality of these trajectories (e.g., diversity) has not been evaluated. Also, some parts of the multi-agent training instances are created using a closed-source LLM (Claude-3.7-Sonnet). LLM generation of synthetic data using frontier (non open-source) models, while widely used (saqib2025teaching; wang2025critique; li2023synthetic; tang2023does), has its limitations in terms of reproducibility. We will release these instances to ensure replicability and comparisons against other methods for training agents. However, the standard issues of using closed-source LLM will make it difficult to compare against other means of generating training data, unless the method is replicated. We will release code and data to support replicability as much as possible.
@@
Since we finetune models on training data with limited scope, models may have learned short-cuts, or overfit to the distribution of the training data. The mixed results on out-of-distribution test data in table 4 and 6 and suggests scope for future work.
@@ Acknowledgements
This work is supported in part by a SUNY-IBM Artificial Intelligence Collaborative Research Alliance grant, an Amazon Research Award 2023, and a NSF NAIRR compute credits grant #240140. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of Amazon.
Appendix A Data Preparation and Trajectory Transformation
A.1 Single-Agent Trajectory Generation
@@
AppWorld appworld-acl24 provides training pairs consisting of input prompts and output states. Specifically, it offers 90 training and 57 development tasks, each with its corresponding output state (i.e., Python solution). We use all 147 tasks to create a fine-tuning dataset for both single-agent and multi-agent setups. To increase the dataset size and introduce solution diversity for each task, we use Llama-3.3-70B-Instruct llama2025instruct33 as the Re-Act agent across 20 different temperature settings. We start at a temperature of $0.05$ and increment by $0.05$ each time, up to $1.0$. To reflect practical scenarios, we first attempt to solve all $147\times 20=2,940$ tasks without providing the output state to the agent. Our base Llama-3.3-70B-Instruct model successfully solves 1,034 out of 2,940 candidate trajectories across the different temperature settings. For the remaining tasks, we provide abstract instructions instead of direct code. We generate these instructions in natural language using Llama-3.3-70B-Instruct. By feeding these abstract instructions to the agent for the 1,906 unsolved candidates, we obtain 1,396 new gold trajectories. For the final 510 candidates, we convert the Python code solutions into pseudocode using Llama-3.3-70B-Instruct. We avoid giving direct Python code as hints because the agent tends to reproduce the code verbatim. Assisting in pseudocode format yields 414 additional gold trajectories. In total, we collect 2,844 gold trajectories. This approach ensures diverse solutions for each task across different temperature settings. When our agent fails to solve tasks in the first phase (i.e., limit exceeds), we use our curated abstract assistance. We did not generate new training tasks for AppWorld, since creating and validating new tasks requires specific environment support, which is challenging to set up.
A.2 Transformation to Multi-Agent Trajectories
We ensure that each agent in the multi-agent system is trained effectively to handle its specific role while maintaining overall system coherence. To achieve this, we have two options at this point to prepare agent-specific training data. One option is direct distillation. First, we use Llama-3.3-70B-Instruct for this purpose. However, open-source models such as Llama-3.3-70B-Instruct are not trained on such long-horizon tasks for multi-agent setting, leading to fragile plans and overall less informative trajectories. On the otherhand, direct distillation of multi-agent trajectories from frontier models is both prohibitively expensive and has fundamental technical limitations. The costs of direct distillation exceed well beyond typical research budgets—our estimate worked out to between 2025,00020-25,000 USD just to obtain the trajectories. To obtain around 3000 trajectories we will have to get around 12000 sample generations from the frontier model. With each generation costing around $2$, this works out to $24,000$ dollars.
In contrast, our cost-efficient trajectory transformation approach leverages existing successful single-agent solutions as a foundation, then systematically decomposes them into coordinated multi-agent workflows. We leverage in-context learning capabilities of LLMs by providing the teacher LLM (Claude-3.7-Sonnet(anthropic2025claude37) with successful single-agent trajectories and prompting it to decompose them into structured multi-agent orchestrations. Given a single-agent trajectory that successfully completes a task in the AppWorld environment, we prompt the LLM to partition the sequence of actions into logical subtasks while preserving the exact execution steps that led to success. Our transformation approach maintains a critical constraint: each executor step (both the reasoning and code) must be identical to the original single-agent trajectory steps. Crucially, we do not allow the model to invent new code or modify existing execution logic - it can only redistribute the original successful steps across subtasks and add exit commands to signal subtask completion. The teacher LLM successfully produces valid 2,708 trajectories from 2844 single agent trajectories. This approach ensures that the generated multi-agent trajectories align with the real execution environment and we further validated these transform actions again by running into the AppWorld environment. This approach proved both economically feasible (cost was only $300$ USD) and effective for our purposes.
While the impact of different distillation strategies can add useful knowledge this is somewhat orthogonal to our key contributions of pareto comparisons of single/multiagent systems, improving multiagents using SLMs through ProST training. The idea of using synthetically distilled dataset is a part of our work but not our main focus. We further analyze error rates for each subtask of the multi-agent trajectories (depict in Figure 8). It shows our dataset covers error correction patterns across all subtasks. We finetune specialized agents of different sizes (7B and 14B), as well as single agents as baselines. Agents in the multi-agent systems are trained on their respective datasets in two ways: (1) standard supervised finetuning, and (2) Progressive Subtask Training (ProST) (details in Section 4.)
A.3 Pipeline Rationale and Verification
The primary reason we used the specific pipeline of Llama-3.3-70B to generate single-agent trajectories and frontier model (i.e., Claude-3.7-Sonnet) to translate into multi-agent trajectories is cost. While our approach enables cost-effective scaling, it also introduces variability. These imperfections can affect fairness by biasing agents toward particular solution styles, and they limit reproducibility since other researchers may obtain different trajectories if different LLMs or prompts are used. To mitigate this, we will release all generated trajectories and prompts, enabling others to train under identical conditions. Understanding how the LLMs structure reasoning in natural language is an interesting study that can shed further light on different reasoning patterns, but one that is beyond the scope and goals of our work.
Notably, both the single-agent and the multi-agent trajectories are fully verified. They are gold trajectories in the sense that using them will solve the corresponding tasks. Both versions share a bulk of their trajectories, with the added parts in multi-agent trajectories being mostly content specific to multi-agent aspects of the problem. Our generated trajectories are fully verified — they are actually validated via execution in the AppWorld environment. We only retain trajectories that actually work for the given task. We conducted further sanity checks on a limited set for prompt engineering to ensure that there are no formatting or structural errors and that the steps remain consistent with the intended task goals. There were no errors in the sub-task goals or the description, although we found some cases where the description can be improved in small ways. The reason for the low-levels of noise here is that the original single agent trajectories we use are gold trajectories – actually are verified. Together, this combination of automatic environment-based validation and manual spot-checking shows that our dataset is both high-quality and task-valid.
We use heterogeneous teachers because generating high-quality action traces and plan-level abstractions requires models with different strengths. This approach mirrors erdogan2025planandactimprovingplanningagents, which similarly separates plan and act supervision, employs open-source models for execution traces, and frontier models for plan generation and expansion. Both approaches use synthetic data generation, annotating ground-truth trajectories with feasible plans to strengthen the planner – a method shown to improve planner quality in. Fairness is further preserved by applying the same evaluation protocols and training pipeline components where applicable across settings, leaving the agent paradigm (single-agent versus multi-agent) and the multi-agent specific learning method (ProST) as the sole intended variables. As reflected by consistent Pareto and TGC/SGC gains across model scales, our observed effect comes from ProST and role specialization, not from differences in teacher models.
Appendix B Implementation Details
B.1 Training and Evaluation Configuration
Our base models are finetuned with Low-Rank Adaptation (LORA) hu2021loralowrankadaptationlarge. We use rank (r=16), alpha ($\alpha$=32) and dropout = 0.05. We use a learning rate of 2.0e-4 with a cosine scheduler. To stabilize training, we set the gradient clipping to 1.0 and the weight decay to 0.01. We train the model for 5 epochs. We use a warm-up ratio of 0.1. The maximum sequence length is set to 20,480 tokens, as our training trajectories are long. All of our experiments are conducted on 4 NVIDIA H100 GPUs with 80GB of memory each. We use Accelerate integrated with DeepSpeed (ZeRO-3) for multi-GPU training. This setup provides efficient memory management and optimization. For both training and evaluation, each agent took  8 GPU hours. For evaluation, we use vLLM kwon2023efficientmemorymanagementlarge servers. For both single-agent and multi-agent setups, we perform inference at T=0.1 and T=0.2 for Test Normal set and Test Challenge, respectively. We set the maximum sequence length to 65536 and enable prefix caching. For the single-agent setting, we allow up to 50 turns for evaluation. For the multi-agent, Orchestrator can plan up to 12 subtasks, and each time Executor allows max 15 turns (for the Test Challenge set, we set max 20 turns) to complete each subtask. A task is considered failed if an agent calls the task completion API with a fail status or doesn’t complete the task within the maximum allotted limits.
B.2 Random Subtask Selection Strategy
B.3 ProST Details
Appendix C Appworld Details
AppWorld appworld-acl24 is a framework designed to assess the ability of autonomous agents in solving real-world problems by interacting with an environment of everyday applications. The framework provides the user with two components: the AppWorld Execution Engine and AppWorld Benchmark.
@@ Appendix C Appworld Details
Model TGC(%) SGC(%)
Qwen-2.5-Coder-7B
SA 7 0.2 0.0
SA 7 [FT] 8.9 2.9
MA 7-7-7 1.7 0.0
MA 7-7-7 [FT] 7.2 3.6
MA 7-7-7 [ProST] 10.5 4.3
Qwen-2.5-Coder-14B
SA 14 4.3 2.2
SA 14 [FT] 16.8 5.8
MA 14-14-14 10.1 3.6
MA 14-14-14 [FT] 13.7 5.0
MA 14-14-14 [ProST] 14.1 5.8
Llama-3.1-8B
SA 8 2.2 0.0
SA 8 [FT] 8.9 2.2
MA 8-8-8 2.2 0.0
MA 8-8-8 [FT] 8.9 4.3
MA 8-8-8 [ProST] 10.8 3.6
Phi-4-14B
SA 14 6.2 2.2
SA 14 [FT] 14.1 3.6
MA 14-14-14 12.2 4.3
MA 14-14-14 [FT] 12.5 7.2
MA 14-14-14 [ProST] 17.8 8.6
Table 6: Performance comparison in single-agent and multi-agent settings on Test Challenge (Test-C) set for Qwen-2.5-Coder-7B, Qwen-2.5-Coder-14B, Llama-3.1-8B, and Phi-4-14B.
Appendix D Extended results
D.1 Generalizability of ProST
The core components of our approach—role specialization, progressive subtask exposure, and Pareto-aware optimization—are not tied to AppWorld’s specific applications. The Orchestrator-Executor-Critic architecture is a generalizable decomposition pattern. Our training strategy, ProST, targets a universal challenge: SLMs’ inability to learn long trajectories under standard fine-tuning. Our ablation studies confirm that the order of subtask introduction matters, suggesting that curriculum design is a critical lever—not just an artifact of AppWorld’s tasks. Furthermore, we successfully apply ProST to non-coding models (Llama-3.1-8B, Phi-4-14B), showing consistent gains across architectures, and observe performance improvements on out-of-distribution tasks, where even modest gains indicate robustness. Using a single, high-quality benchmark for comprehensive evaluation is standard practice for complex problems due to a range of reasons. Firstly, working on complex task domains often requires significant engineering effort and environment setup cost. Many prominent works in software engineering and agent evaluation use single long horizon domains for this reason. For instance, the widely-cited SWE-bench jimenez2024swebench paper has become the gold standard for evaluating coding agents and notable works (yang2024sweagentagentcomputerinterfacesenable; zhang2024autocoderoverautonomousprogramimprovement; ma2024lingmaswegptopendevelopmentprocesscentric; pan2025trainingsoftwareengineeringagents; golubev2024search; xia2024agentlessdemystifyingllmbasedsoftware; xie2025swefixertrainingopensourcellms) attempt to resolve programming tasks by interactive coding agents. Making progress on AppWorld represents useful advances in the complex problem space. The complex tasks in AppWorld are defined over every-day apps in the AppWorld benchmark, which is comparable to SWE-bench in scope and difficulty (see AppWorld details in Appendix C). This is a challenging benchmark and there are other recent works that focus exclusively on improving performance on this benchmark as well (chen2025reinforcementlearninglonghorizoninteractive; gupta2025leveraging). A recent multi-agent analysis paper also showed that AppWorld tasks pose substantial reasoning and coding challenges cemri2025multiagentllmsystemsfail. Another recent work qi2025webrltrainingllmweb evaluate solely their method on a long-horizon Web benchmark WebArena due to the challenges of complexity of configuring a second long horizon benchmark. Another limiting factor in exploring other complex tasks benchmarks is scacity of training data for multi agent systems. We spent a significant amount of time creating multi step reasoning training data (i.e for AppWorld) following rigorous processes. Thus, we argue that applying any other long-horizon benchmark’s tasks to the ProST training paradigm would also yield significant performance improvements compared to standard fine-tuning.
Appendix E Prompts Details